# Rewarding Curse: Analyze and Mitigate Reward Modeling Issues for LLM Reasoning

### v1

@@ Abstract
Inference-time scaling techniques enable large language models (LLMs) to solve complex reasoning problems, in which the reward model (RM) plays an important role in selecting correct paths. In this paper, we present a systematic analysis of the RM’s performance in downstream reasoning tasks, which are overlooked in related works. We explore the discriminative abilities of RM from three perspectives and draw some meaningful conclusions: (1) Question difficulty: RM can impair LLM’s performance when solving simple problems; (2) Sampling number: RM struggles to distinguish between correct and incorrect responses as sampling increases and encounters low-frequency negatives; (3) Search diversity: RM requires a response distribution with moderate diversity to achieve optimal performance. Based on these findings, we design a new algorithm called Optimal Clustering Tree Search (OCTS) to mitigate these RM’s issues in the inference phase simultaneously. Experimental results demonstrate our method effectively enhances LLM’s reasoning capabilities, achieving up to a 3.2% improvement in accuracy.
Rewarding Curse: Analyze and Mitigate Reward Modeling
Issues for LLM Reasoning
Jiachun Li1,2, Pengfei Cao1,2, Yubo Chen1,2 Jiexin Xu3, Huaijun Li3, Xiaojian Jiang3, Kang Liu1,2, Jun Zhao1,2 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3China Merchants Bank {jiachun.li, pengfei.cao, yubo.chen, kliu, jzhao}@nlpr.ia.ac.cn
@@ 1 Introduction
The remarkable achievements of OpenAI’s o1 have sparked a wave of research into inference-time scaling techniques in reasoning tasks (OpenAI, 2024; DeepSeek-AI et al., 2025; Zeng et al., 2024). In these works, the reward model (RM) plays a pivotal role in guiding large language models (LLMs) to search for the correct reasoning path in the inference phase (Wang et al., 2024b; Setlur et al., 2024; Zhang et al., 2024). When combined with strategies like Best-of-N (BoN) or Monte Carlo Tree Search (MCTS), the RM can recall the best response based on the LLM’s thorough exploration of the solution space, thereby improving the reasoning performance.
Due to the significance of reward models, a series of works have been proposed to study their performances, which provide mixed results (Lambert et al., 2024; Liu et al., 2024b; Zheng et al., 2024). On one hand, some works show that RM combined with advanced search strategies (e.g. MCTS) can effectively enhance the LLM’s reasoning capabilities (Setlur et al., 2024; Jiang et al., 2024; Wan et al., 2024). On the other hand, some works like DeepSeek-R1 point out that neural reward models suffer from reward hacking, which limits their effectiveness in enhancing LLM reasoning capabilities (DeepSeek-AI et al., 2024, 2025). These conflicting findings motivate the need for a systematic analysis of the reward model’s performance. However, related works mainly focus on evaluating the overall performance of the RM, without delving into how various factors influence these models in downstream reasoning tasks (Wang et al., 2024b; Zhang et al., 2024).
@@
In this work, we focus on investigating the factors that influence the RM’s performance in math reasoning tasks and analyzing its existing issues. Specifically, we begin by mathematically modeling the RM-based inference process to identify the key factors in it, including questions, sampling number, and search parameters. Then, we conduct experiments centered around these factors to analyze their impact on the RM’s performance: (1) For the questions, we test the performance of BoN and MCTS across different question difficulty levels, demonstrating that the introduction of the RM primarily enhances the model’s reasoning ability on difficult problems, while it can impair the performance on simpler problems. (2) For the sampling number, we track the RM’s discriminative ability under different numbers and find that more samples make it increasingly difficult for the model to distinguish between positive and negative responses. Through statistical analysis, we attribute this issue to the RM’s inverse long-tail phenomenon, where they tend to assign higher scores to incorrect responses that occur with low frequency. (3) For the search parameters, we primarily study the impact of certain parameters that control search diversity, including sampling temperature, tree structure, etc. We find that RM requires a response distribution with moderate diversity to achieve optimal performance. Excessive diversity generates high-quality negative examples, which lead to performance degradation.
Based on these findings, we design a novel RM-based inference algorithm called Optimal Clustering Tree Search (OCTS), which consists of three steps: exploration, selection and expansion. In exploration, we directly generate multiple responses like BoN, stopping the algorithm and returning the majority-voted answer if the question difficulty is low. This step mitigates the performance drop caused by the RM on simple problems. In selection, we cluster responses based on their answers and select the current optimal path by considering both frequency and reward scores, which mitigates the inverse long-tail issue. In expansion, we use the early steps of the selected path as prefixes and directly generate all of the remaining steps as new paths. This reduces the intermediate states of the tree, alleviating the performance loss caused by excessive diversity. We conduct extensive experiments to compare our method with other baselines. The results not only indicate our method is effective in improving RM-based reasoning abilities but also demonstrate the rationality of our former findings.
Our main contributions are as follows: (1) We systematically analyze the key factors influencing RM’s performance in downstream reasoning tasks under the guidance of mathematical modeling. (2) Through comprehensive experiments, we have three meaningful findings, including RM being harmful when facing simple questions, RM struggling to distinguish low-frequency negative samples, and RM performing worse on high-diversity distributions. (3) We propose the OCTS algorithm, which effectively enhances the RM-based reasoning performance of LLMs. Experimental results demonstrate that we can improve up to 3.2% accuracy compared to other baselines.
2 Preliminaries
In this section, we first evaluate the performance of various RMs on the math reasoning task as a foundation for our subsequent analysis (§2.1). Then, we mathematically model the RM-based inference process to help identify the key factors we study in this work (§2.2).
2.1 Overall Performance Experiment
@@
A commonly used method to evaluate reward models is BoN, which generates multiple responses and selects the best one based on the reward score. The settings in this experiment are as follows:
Baselines
In addition to BoN, we also set two baselines: SC and Oracle. For SC, we sample $N$ responses and select the major voting answer. For Oracle, we assume the existence of an oracle selector, which directly selects the correct answer from the $N$ samples generated. We regard SC as a measure of the LLM’s own reasoning capabilities, with Oracle serving as the ceiling of the performance.
Models
For the policy model, we select some representative open-source models: Mistral-7B (Jiang et al., 2023), Gemma2-9B (Rivière et al., 2024a), Llama2-13B (Touvron et al., 2023), Llama3.1-8B (Rivière et al., 2024b), Qwen2.5-3B and Qwen2.5-14B (Yang et al., 2024). For the reward model, we evaluate some advanced models: Shepherd-Mistral-7B-PRM (Wang et al., 2024b), ArmoRM-Llama3-8B (Wang et al., 2024a), Skywork-Llama-3.1-8B (Liu et al., 2024a), Skywork-o1-PRM-Qwen-2.5-7B (o1 Team, 2024). These models encompass a variety of base models and demonstrate commendable performance (see Appendix A for details).
@@
Datasets
Following previous works (Snell et al., 2024; Brown et al., 2024; Qi et al., 2024), we select MATH-500 (Hendrycks et al., 2021; Lightman et al., 2024) as the dataset for our experiments, which consists of high-school competition-level math problems. Unless otherwise specified, all experiments in this paper are conducted on it.
@@
Figure 1 shows the main results of the evaluation (more results, including more datasets and the MCTS in Appendix B). We can conclude that: Advanced reward models have limited performance on the downstream math reasoning task. For most LLMs (except for Gemma2-9B), BoN only provides minor improvements over SC (<5%). On Qwen-2.5-3B, the BoN for all reward models exhibits lower accuracy than SC, indicating that the RM can even undermine the model’s inherent reasoning performance. Besides, Oracle significantly outpaces other baselines, suggesting that the performance bottleneck lies in the RM’s discriminative ability rather than the LLM’s generative capability. Therefore, investigating factors that influence the RM’s performance and analyzing existing issues are crucial for enhancing LLM’s reasoning performance, which is also the focal point of this paper.
@@ 2.2 Mathematical Modeling
During the inference phase, the first step is to input the question $q$ and generate multiple responses $\mathbb{R}$:
$\mathbb{R}=S(M(q),N;\Phi)$ (1)
where $M(q)$ denotes the output distribution of the policy model after inputting the question, $N$ denotes the number of samples and $\Phi$ denotes the parameters of the search strategy $S$ (such as sampling temperature). After that, we use a scoring function $f$ to select the best response $\hat{r}$ from $\mathbb{R}$:
$\hat{r}=\underset{r\in\mathbb{R}}{\arg\max}\,f(r)$ (2)
There are various ways to define $f$. Since we focus on studying the RM’s performance, we set $f$ as the score output by the reward model.
In this work, we primarily investigate factors that affect the performance of the reward model (i.e. $f$ in Eq. 2). To this end, we modify the variables in Eq.1 to observe the accuracy of predicted $\hat{r}$ under different $\mathbb{R}$. Specifically, we conduct experiments on two representative search strategies $S$ (i.e. BoN and MCTS), analyzing from three perspectives: the question $q$, the sampling number $N$, and the search parameters $\Phi$.
3 Question Difficulty: RM is harmful when facing simple questions
In this section, we primarily investigate how different questions affect the reward model’s performance. To answer this question, we first model different questions using difficulty as a metric (§3.1). Then, we conduct experiments on the BoN (§3.2) and MCTS (§3.3) paradigms to analyze the influence of question difficulty.
3.1 Question Difficulty Modeling
Following the approach of former works, we use question difficulty as a metric to classify different questions (Lightman et al., 2024; Snell et al., 2024). Specifically, we bin the policy model’s pass@1 rate (estimated from 10 samples) on each question into five quantiles, each corresponding to increasing difficulty levels. For example, if the model answers correctly only 0 or 1 time, the question is classified as the hardest level 5. Conversely, if the model answers correctly more than 8 times, the question is classified as the easiest level 1. Besides, we also study the difficulty approximation without the ground truth and report results in Appendix C.
3.2 BoN Performance across Difficulty
Experimental Setup
We employ the BoN strategy to evaluate the RM’s reasoning performance across varying difficulty levels. Based on results in Figure 1, we select the best-performing Skywork and Skywork-o1 as the ORM (Outcome Reward Model) and PRM (Process Reward Model) for our subsequent experiments. We sample 32 examples from Qwen2.5-3B and Llama3.1-8B models for each question and compare the accuracy.  [^1] Unless otherwise specified, all subsequent experiments utilize Qwen2.5-3B as the primary policy model and Llama3.1-8B as the supplementary model.
@@ 3.2 BoN Performance across Difficulty
Main Results
We illustrate the experimental results in Figure 2, from which we conclude that: Compared to SC, BoN performs worse on simple questions but better on difficult questions. From the easiest level 1 to the hardest level 5, the performance of SC gradually declines, while BoN’s accuracy transitions from lagging behind SC to surpassing it.
Experiments on More Datasets
To validate the generalizability of our conclusion, we also repeat the experiment on two more math reasoning benchmarks: GSM8K (Cobbe et al., 2021) and OlympiadBench (He et al., 2024). We present the details and results of this experiment in Appendix D, which further confirm our above conclusion.
3.3 MCTS Performace across Difficulty
Can more advanced search strategies help to mitigate the above performance decrease on simple questions? To study this problem, we use the MCTS as the search strategy and compare performance across different difficulty levels.
Experimental Setup
In MCTS, we use two different scoring functions $f$ to select the final response: MCTS-SC and MCTS-RM (more functions in Appendix B). For the former, we employ a majority voting method to select the final answer. For the latter, we choose the path with the highest reward score as the final answer. We use ORM and PRM to evaluate the Q-value of final states and perform 32 rollouts over 200 questions.
(a) ORM
(b) PRM
@@ 3.3 MCTS Performace across Difficulty
Figure 3 shows the main results of this experiment. We can observe that, although MCTS provides a certain improvement over BoN, the accuracy of MCTS-RM still lags behind that of SC for low-difficulty problems (see level 1 and 2 in Figure 3(a)). In contrast, MCTS-SC achieves higher accuracy on easy questions but performs worse on harder questions compared to MCTS-RM. This indicates that: (Cl.1) The introduction of the RM can hinder the LLM’s performance on simple problems but improve its effectiveness when addressing complex tasks. This pattern is not limited to specific search strategies.
4 Sampling Numbers: RM struggles to distinguish low-frequency negatives
In this section, we study the impact of the sampling numbers on the RM’s performance. To this end, we first investigate the overall performance of BoN and MCTS as the sampling numbers scales (§4.1). Next, we evaluate the reward model’s capability to differentiate between correct and incorrect responses under different $N$4.2). Finally, we conduct a statistical analysis of the negative samples to explain the reasons behind our findings (§4.3).
4.1 Overall Performance
Experimental Setup
Recent works (Brown et al., 2024) demonstrate the LLM’s coverage of correct answers (i.e. the Oracle in §2.1) increases as the sampling number grows, whereas the accuracy does not fully scale with $N$. Based on it, we further investigate whether introducing better RMs and the MCTS strategy can reduce the gap between coverage and accuracy. Specifically, for BoN, we vary $N$ from 1 to 128, while for MCTS, we set $N$ from 1 to 64. For the implementaion of MCTS, we utilize the RM’s score as the scoring function $f$ (i.e. MCTS-RM in §3.3), which will serve as the default in the subsequent experiments.
(a) BoN
(b) MCTS
Figure 4: Two inference methods performance across difference sampling numbers.
Main Results
The changes in accuracy and coverage are shown in Figure 4. From the results, we can conclude that: (1) The performance of BoN inference can not be scaled with $N$. For both PRM and ORM, the accuracy of BoN inference plateaus beyond approximately 50 samples in Figure 7(a). As $N$ increases, the gap between their accuracy and coverage becomes progressively larger. (2) The accuracy of MCTS also cannot scale continuously with the sampling number. As shown in Figure 7(b), the accuracy of MCTS also plateaus before reaching a relatively small $N$. In contrast, the Oracle setting consistently increases, leading to a persistently widening gap between accuracy and coverage.
4.2 Discriminative Experiments on RM
In the context of increasing coverage, whether the accuracy improves or not is predominantly contingent upon the discriminative capacity of the reward model. Thus, we design experiments to observe the performance of the RM in distinguishing positive and negative instances as $N$ increases.
(a) BoN
(b) MCTS
Case Analysis

### v2

# Fixing the Broken Compass: Diagnosing and Improving Inference-Time Reward Modeling

@@ Abstract
Inference-time scaling techniques have shown promise in enhancing the reasoning capabilities of large language models (LLMs). While recent research has primarily focused on training-time optimization, our work highlights inference-time reward model (RM)-based reasoning as a critical yet overlooked avenue. In this paper, we conduct a systematic analysis of RM behavior across downstream reasoning tasks, revealing three key limitations: (1) RM can impair performance on simple questions, (2) its discriminative ability declines with increased sampling, and (3) high search diversity undermines RM performance. To address these issues, we propose CRISP (Clustered Reward Integration with Stepwise Prefixing), a novel inference-time algorithm that clusters generated reasoning paths by final answers, aggregates reward signals at the cluster level, and adaptively updates prefix prompts to guide generation. Experimental results demonstrate that CRISP significantly enhances LLM reasoning performance, achieving up to 5% accuracy improvement over other RM-based inference methods and an average of 10% gain over advanced reasoning models.
@@ 1 Introduction
The remarkable achievements of OpenAI’s o1 have sparked a wave of research into inference-time scaling techniques in reasoning tasks (OpenAI, 2024; DeepSeek-AI et al., 2025; Zeng et al., 2024). Some works aim to enhance models during the training phase, employing reinforcement learning (RL) (Xie et al., 2025; Qu et al., 2025) or supervised fine-tuning (SFT) (Ye et al., 2025; Muennighoff et al., 2025) on high-quality data to equip models with the ability to generate long chains of thought (CoT). Other approaches focus on inference-time optimization, using reward model (RM)-based search strategies such as Monte Carlo Tree Search (MCTS) to guide the model toward more efficient solution paths (Wang et al., 2024b; Setlur et al., 2024; Zhang et al., 2024).
Driven by the great success of the DeepSeek-R1 series (DeepSeek-AI et al., 2025), recent efforts have predominantly focused on reproducing its performance from a training-centric perspective (Muennighoff et al., 2025; Ye et al., 2025; Xie et al., 2025), while largely overlooking inference optimization methods. Although R1-style works achieve strong performance on tasks such as math reasoning, they have been shown to suffer from serious issues such as overthinking (Chen et al., 2024; Sui et al., 2025) and limited task generalization (Zhang et al., 2025a; Zheng et al., 2025). These issues, however, can be mitigated through RM-based inference techniques. For example, on the commonsense reasoning task CSQA (Talmor et al., 2019), DeepSeek-R1-7B (DeepSeek-AI et al., 2025) achieves 64.8 accuracy with an average of 3,613 tokens. In contrast, our RM-based inference method, applied to its base model Qwen2.5-Math-7B (Yang et al., 2024b), reaches a higher accuracy of 72.0 using only 1,100 tokens on average. Therefore, optimizing inference-time reasoning remains a critical direction, particularly for smaller models.
How can we further improve the reasoning performance of LLMs at inference time? Revisiting R1-style work, one key insight is their identification of the reward hacking issue during RL training, which they address using rule-based reward functions, ultimately improving performances (Liu et al., 2024b; DeepSeek-AI et al., 2025; Gao et al., 2023). This raises a natural question: Can we similarly analyze the issues of the reward model at inference time and mitigate them to enhance the LLM’s reasoning ability?
@@
In this work, we investigate the factors affecting reward model performance at inference time and propose methods to mitigate its limitations. Specifically, we begin by mathematically modeling the RM-based inference process to identify its key influencing factors: the input questions, the number of sampled responses, and the search parameters. Then, we conduct targeted experiments to analyze the impact of each factor on RM performance: (1) Input question: We test the performance of BoN and MCTS across different question difficulty levels and demonstrate that RM-based inference significantly impairs performance on simple questions. (2) Sampling number: We analyze the RM’s discriminative ability under different numbers $n$ and observe that its performance deteriorates as $n$ increases. The statistical analysis attributes this degradation to an inverse long-tail phenomenon, wherein the RM tends to assign higher scores to low-frequency, incorrect responses. (3) Search parameters: We focus on parameters controlling search diversity, such as sampling temperature and MCTS tree structure. Our results show that RM performs best under moderate diversity, while excessive diversity undermines reasoning accuracy.
To mitigate the former issues in RM-based inference, we design a novel algorithm called CRISP (Clustered Reward Integration with Stepwise Prefixing). CRISP operates in an iterative fashion, where each round begins by sampling reasoning paths conditioned on a dynamic prefix set. These paths are then clustered by their final answers, allowing the algorithm to aggregate reward signals at the cluster level and thereby attenuate the RM’s tendency to mis-rank rare but incorrect outputs. We further incorporate an early termination mechanism based on cluster cardinality, which enables efficient inference on simple questions and alleviates RM instability in such cases. Finally, high-scoring paths from dominant clusters inform the construction of stepwise prefixes for the next sampling round, enabling tighter control over search diversity by limiting the number of intermediate states explored. We conduct extensive experiments to compare our method with other baselines. The results not only indicate that our method is effective in improving RM-based reasoning abilities, with accuracy gains of up to 5%, but also validate the soundness of our earlier findings. Moreover, compared to DeepSeek-R1 models of the same scale, our method reduces average token usage by up to 90%, while achieving an average accuracy improvement of 10% on non-mathematical tasks.
Our main contributions are as follows: (1) We draw three critical findings based on a systematic analysis of RM behavior during inference: RM degrades performance on simple questions, fails to effectively distinguish low-frequency incorrect samples, and performs suboptimally under excessive search diversity. (2) We propose CRISP, a novel inference-time algorithm that clusters generated reasoning paths by final answers, aggregates reward signals at the cluster level, and adaptively updates prefix prompts to guide generation, effectively mitigating the shortcomings of reward models at inference time. (3) Extensive experiments demonstrate that CRISP consistently outperforms both inference-time and training-time baselines, with accuracy improvements of up to 5% compared to other RM-based inference methods, and an average of 10% over R1 models in non-mathematical reasoning tasks. The code is available at https://github.com/BugMakerzzz/CRISP.
2 Overall Performance of Reward Models in Inference-Time
@@
In this section, we evaluate the inference-time performance of the current reward model as a preliminary experiment. Specifically, we compare the accuracy of Best-of-N (BoN), which generates multiple responses and selects the best one based on the reward score.
Experimental Setup
For the policy model, we select representative open-source models: Gemma2-9B (Rivière et al., 2024a), Llama3.1-8B (Rivière et al., 2024b), Qwen2.5-3B and Qwen2.5-14B (Yang et al., 2024a). For the reward models, we select two outcome reward models (ORMs): ArmoRM (Wang et al., 2024a) and Skywork-Llama-3.1-8B (Liu et al., 2024a), and two process reward models (PRMs): Shepherd-Mistral-7B-PRM (Wang et al., 2024b) and Skywork-o1-PRM-Qwen-2.5-7B (Team, 2024). These models demonstrate commendable performance on related benchmarks (see Appendix C for details). As for the evaluation data, following previous works (Snell et al., 2024; Brown et al., 2024; Qi et al., 2024), we select MATH-500 (Hendrycks et al., 2021; Lightman et al., 2024), which consists of high-school competition-level math problems. In addition to BoN, we also set two baselines: SC and Oracle. For the former, we select the major voting answer from $n$ responses. For the latter, we directly recall the existing correct answer from the generated samples, which serves as the performance ceiling.
@@
Figure 1 shows the main results of the evaluation (see Appendix D for more results). We can conclude that: Advanced reward models have limited performance on the downstream math reasoning task. For most LLMs, BoN only provides minor improvements over SC ($5\%$). Specifically, on Qwen2.5-3B, the BoN for all reward models exhibits lower accuracy than SC, indicating that the BoN inference method has limited reasoning performance. Besides, Oracle significantly outpaces other baselines, suggesting that the performance bottleneck lies in the RM’s discriminative ability rather than the LLM’s generative capability. Therefore, identifying and mitigating the factors that impair the RM’s performance during inference are crucial for enhancing LLM’s reasoning ability.
3 Probing RM-based Inference Issues
@@ 3.1 Mathematical Modeling
During the inference phase, the first step is to input the question $q$ and generate multiple responses $\mathcal{R}$:
$\mathcal{R}=\mathcal{S}(\mathcal{M}(q),n;\Phi)$ (1)
where $\mathcal{M}(q)$ denotes the output distribution of the policy model after inputting the question, $n$ denotes the number of samples and $\Phi$ denotes the parameters of the search strategy $\mathcal{S}$ (such as sampling temperature). After that, we use a scoring function $f$ to select the best response $\hat{r}$ from $\mathcal{R}$:
$\hat{r}=\underset{r\in\mathcal{R}}{\arg\max}\,f(r)$ (2)
To analyze the performance of the reward model, we define $f$ as the score predicted by the RM. Our work focuses on identifying key factors that influence RM performance. To this end, we vary the components in Eq.1 to observe the accuracy of predicted $\hat{r}$ under different $\mathcal{R}$. Specifically, we study three main factors through probing experiments: the input question $q$, the sampling number $n$, and the search parameters $\Phi$.
3.2 Experimental Setup
For reward models, based on results in Figure 1, we select the best-performing Skywork and Skywork-o1 as the ORM and PRM for our subsequent experiments. Regarding policy models, we use Qwen2.5-3B and Llama3.1-8B throughout our experiments. To ensure that our findings are not specific to a particular strategy, we conduct all experiments using both BoN and MCTS. As for evaluation data, we employ the MATH-500 dataset in our main text, and provide additional results on GSM8K (Cobbe et al., 2021) and OlympiadBench (He et al., 2024) in the appendix.
3.3 Input Question: Reward Model Underperforms on Easy Questions
Question Difficulty Modeling
We first investigate how different questions affect the RM’s performance. Following former works, we use question difficulty as a metric to classify different questions (Lightman et al., 2024; Snell et al., 2024). We bin the policy model’s pass@1 rate (estimated from 10 samples) on each question into five quantiles, each corresponding to increasing difficulty levels. For example, If the model answers correctly 0 or 1 time, the question is level 5 (hardest). If it answers correctly more than 8 times, the question is level 1 (easiest). To facilitate a holistic and rigorous evaluation of the problem difficulty, we present results based on dataset difficulty partitions in Appendix N. We also include experiments demonstrating difficulty estimation in the absence of ground truth answers in Appendix E.
BoN Performance
After categorizing the data by difficulty, we analyze the BoN performance across different levels. We sample 32 examples from each question and illustrate the accuracy in Figure 3, from which we can conclude that: Compared to SC, BoN performs worse on simple but better on difficult questions. From the easiest level 1 to the hardest level 5, the accuracy of SC gradually declines, while BoN transitions from lagging behind SC to surpassing it. We also repeat the experiment on two more math reasoning benchmarks and present the results in Appendix F, further confirming our conclusion.
@@ 3.3 Input Question: Reward Model Underperforms on Easy Questions
(c) ORM
(d) PRM
@@
In MCTS, we use two different scoring functions $f$ to select the final response for comparison: MCTS-SC and MCTS-RM (more functions in Appendix D). For the former, we employ a majority voting method for selection. For the latter, we choose the path with the highest reward score. We perform 32 rollouts over 200 questions, demonstrating the results in Figure 3. Although MCTS provides improvement over BoN, the accuracy of MCTS-RM still lags behind that of SC for low-difficulty problems (see levels 1 and 2 in Figure 3). Besides, MCTS-SC achieves higher accuracy on easy questions but performs worse on harder questions compared to MCTS-RM. These indicate that: (Cl.1) The introduction of the RM can hinder the LLM’s reasoning performance on simple problems. This pattern is not limited to specific inference strategies.
3.4 Sampling Number: RM struggles to distinguish low-frequency negatives